# Early warning indicators via latent stochastic dynamical systems

### 30 Dec 2023 v2

@@ 3 Experiments
We now apply the method outlined in the previous sections to datasets collected from patients with epileptic seizures. Our objective is to analyze the latent dynamics and illustrate our early warning strategy.
@@ 3.1 Data Preprocessing
Electroencephalogram (EEG) entails the capture of signals using electrodes positioned on the scalp, thereby recording the electrical activity during epileptic seizures. The EEG dataset encompasses both pre-ictal and ictal data, organized in a matrix format of channels over time, with a sampling rate of 256 points per second. To facilitate analysis, each electrode signal undergoes rescaling to an interval of $[-0.5,0.5]$ and is subsequently subsampled by averaging every sixteen points, resulting in the formation of the dataset. In Fig. 2, there is a sample image that the doctor labels $Time=500$ as the separation time for pre-ictal and ictal period. As different doctors may assign varying labels over time, here $Time=500$ is utilized as a reference.
@@ 3.2 Dimension Reduction with Directed Anisotropic Diffusion Map
To construct the diffusion matrix, the data from all electrodes at each time point is viewed as a high-dimensional node in the diffusive graph. Notably, electrode signals display abnormal fluctuations preceding the onset of an epileptic seizure. Timely detection of abnormalities in electrical signals holds the potential to provide early warnings for epileptic seizures. Such early warnings can enable specific intervention measures aimed at mitigating or avoiding epileptic seizures. We suggest identifying early warnings within a lower-dimensional space through the application of directed anisotropic diffusion maps.
3.2.1 Diffusion Matrix and Spectral Gap
@@
The diffusion matrix is constructed by the diffusion kernel and shows the transition among all high-dimensional nodes in the diffusive graph, which exhibits the transition probability from one point to another. The directed drift term $\nabla f(x)=|\dot{x}|$ is data-driven, where
$\dot{x}=\frac{1}{2}\left(\frac{x(t_{i+1})-x(t_{i})}{t_{i+1}-t_{i}}+\frac{x(t_{i})-x(t_{i-1})}{t_{i}-t_{i-1}}\right).$
Then we apply the directed diffusion to the preprocessed dataset. The drift term can be thought as an enhanced force.
Figure 3: Comparison of diffusion map and directed anisotropic diffusion map. The top row presents the kernel matrices, and the bottom row represents the eigenvalues. For all cases, the plots we show here are under $\epsilon=1$: left column ((a)(c)) is constructed when $\nabla f=0$, while the right column ((b)(d)) is calculated when $\nabla f(x)=|\dot{x}|$.
@@
As shown in the top row of Fig. 3, most points have a measure greater than $0.5$ with another point in $(a)$, indicating that the transition probability to other points is equivalent. In $(b)$, the data points within the red frame have higher transition probabilities to each other, while the transition probabilities to the points outside gradually decrease. Through directed diffusion, the transition probability of sample points inside the red box to outside the red box is reduced. Correspondingly, the transition probability from the sample point outside the red box to the data point inside the red box also decreases accordingly. In this way, the diffusion activity intensity between the points inside the red box and the points outside the red box in the original high-dimensional time series is forced to attenuate, so the two regions will be better separated.
The diffusion matrix has a discrete sequence of eigenvalues $\{\lambda_{i}\}_{i\geq 1}$ and eigenvectors $\{\phi_{i}\}_{i\geq 1}$ such that $1\geq\lambda_{1}\geq\lambda_{2}\geq\cdots$. The eigenvalues affect diffusive strength, and the spectral gap is an indicator to help us find the effective dimension of latent space. In the bottom row of Fig. 3, the first five eigenvalues of the diffusion map have similar gaps, while the first eigenvalue of the directed anisotropic diffusion map shows a distinct spectral gap from the second one. Then the largest eigenvalue is separated from the others.
3.2.2 Low-dimensional Embedding over Time
@@
Figure 4: Comparison of low-dimensional embedding patterns with (a) diffusion map ($\nabla f=0$) and (b) directed anisotropic diffusion map ($\nabla f=|\dot{x}|$).
The directed anisotropic diffusion map discovers a latent coordinate denoted as $\Phi_{1}$ by the gap between two adjacent eigenvalues, as shown in Fig. 3 bottom row. This suggests that the first diffusion coordinate is enough to provide an economical representation of the original EEG dataset. We check the interpretability of the data-driven observations by displaying the coordinates in Fig. 4. The diffusion map projects data points to divergent locations. Moreover, the anisotropic diffusion map partitions the dataset into two states with a distribution shift in the latent space, expended by the first leading diffusion coordinate. From $(a)$ in Fig. 4, the data points are fluctuant all the time with the same expectation. However, in $(b)$, the drift term induced by the gradient of data $\nabla f(x)$ helps to differentiate two groups of points at $Time\approx 600$ with different expectations.
@@ 3.3 Latent Stochastic Dynamical Systems
An interesting aspect of our research involves the dynamic behaviors exhibited by latent dynamical systems. In this section, we aim to elucidate the connection between the latent stochastic differential equation and the original high-dimensional time evolutionary data.
Following dimension reduction through the directed anisotropic diffusion map, we obtain the projected dataset onto the first diffusion coordinate $\Phi_{1}$. The projected data plotted over time displays the transition between two states: pre-ictal state and ictal state. Therefore, we assume that the stochastic differential equation in the latent space has two meta-stable states and rescale the data to an interval of $[-1,1]$ for convenience. Given data in Fig. 4 (b), we would like to obtain the latent stochastic differential equation in the form of Eq. (4). To get meaningful representations of drift and diffusion coefficients, we assume the utilization of a polynomial basis for approximating both $\mu(z)$ and $\sigma(z)$ and employ neural networks to approximate the latent equations. The inputs of networks are the basis functions of the coordinates and the time step is $0.0625$. The latent space contains $1500$ time points and $1499$ pairs of $\{(z(t_{i+1}),z(t_{i}))\}$ are available. We randomly select $1200$ pairs to train the networks and validate the identification accuracy on the rest $299$ pairs.
To find the explicit form of drift and diffusion coefficients in the unknown stochastic differential equations, all the unknown drift and diffusion terms are parameterized with neural networks, with a polynomial basis as the input vector. Then using the loglikelihood method in Sec. 2.2, the learnt stochastic differential equation under the coordinate $\Phi_{1}$ has the form
$dz(t)=\mu_{1}(z(t))dt+\sigma_{1}(z(t))dB_{t},$ (9)
where $\mu_{1}(z)=0.0219+0.7128z-0.0774z^{2}-1.1677z^{3}$, and $\eta_{1}(z)=-0.5115$.
Moreover, we adopt density as the validation metric for the testing data (consisting of $299$ pairs). In Fig. 5, we compare the density of the estimated $z(t_{i+1})$ obtained through the approximation of the latent stochastic differential equation (Eq. (9)) with that of the ground truth from corresponding data points in Fig.4(b). The alignment between our approximation and the true data density is evident. Additionally, the presence of two peaks in the density plot signifies the existence of two meta-stable states.
Figure 5: The validation of the neural network approximation with regard to $\Phi_{1}$.
@@ 3.4 Early Warning Indicators
For $\Phi_{1}$, we employ a time series trace comprising $1500$ data points to calculate the indicator over time and compare it with the standard deviation. The time window $l$ is set to $100$, moving forward in each time unit. The Onsager-Machlup indicator is computed in each time window using the Eq. (6). When applying the Eq. (7) to compute the sample entropy, the parameters are kept constant at $m=10$, $p=1$, $q=4$, and $r=0.5$. The values of the two indicators over time are depicted in red in Figure 6 and the standard deviation is in green.
Assume that a transition will occur if the indicator surpasses a predefined threshold. Two patterns are observed in the coordinate signal over time. Until approximately $600$, the Onsager-Machlup indicator remains negative, and the sample entropy consistently equals 00. In contrast, the standard deviation, commonly employed as an indicator, undergoes a change at $Time\approx 200$. This could lead to erroneous judgments. This implies that both the Onsager-Machlup indicator and the sample entropy indicator offer more precise warning points compared to the standard deviation. Consequently, this identified change point can assist doctors in distinguishing between pre-ictal and ictal periods, facilitating automated labeling processes.
Figure 6: The Onsager-Machlup indicator $(a)$ and the sample entropy $(b)$ of raw data projected onto coordinate $\Phi_{1}$, compared with the standard deviation(green).
@@
Assigning the width of each time window as $100$, Fig.7 shows the transition probability. The first $100$ points are selected as starting points, generating $100$ samples. The split point is set as $0.8$, delineating regions $A=[0.8,\infty)$ and $B=(-\infty,0.8)$. In Fig. 7, we calculate the transition probability by Eq. (8) of state for each time step $t$ starting from the initial state. Assuming a transition occurs when the probability surpasses $0.5$, we designate the point with a probability equal to $0.5$ as the warning point. AOver time, distinct patterns emerge in the signal of the first coordinate, with warning point occurring around $Time=600$ for $\Phi_{1}$.
Figure 7: The transition probability of raw data projected onto coordinate $\Phi_{1}$. The red dot indicates the transition probability equal to $0.5$.
@@ 4 Conclusion
Predicting epileptic seizures from brain signals is inherently challenging and time-consuming, primarily due to the high dimensionality and complexity of the data. In this paper, we propose an efficient method for automatically identifying abnormalities in EEG signals. Our approach leverages directed anisotropic diffusion map for dimensionality reduction and establishes a stochastic dynamical system model in the low-dimensional latent space. This enables effective and time-saving detection of early warning signals associated with epilepsy.
@@
Early warning of epileptic seizures is of paramount importance for individuals with epilepsy. The abrupt change is caught for early warning in the latent space, where normal state and ictal state can be viewed as two meta-stable states. The ability to identify transitions between meta-stable states plays a pivotal role in predicting and controlling brain behavior. We employ stochastic differential equations to model the nonlinear dynamics in the latent space. Furthermore, we derive three effective warning signals—namely, the Onsager-Machlup Indicator, Sample Entropy Indicator, and Transition Probability Indicator—utilizing information from the latent coordinates and the latent stochastic dynamical systems. These indicators enhance the robustness and accuracy of early warning systems for epileptic seizures. This framework of learning latent stochastic systems and detecting abnormal dynamics has the potential to extend to general scenarios for other complex high dimensional time evolutionary data.

### 05 Apr 2024 v3

@@ 3 Experiments
We now apply our method to the real datasets collected from epileptic patients. Our objective is to analyze the latent dynamics and illustrate our early warning indicators.
@@ 3.1 Data Preprocessing
Electroencephalogram (EEG) captures signals from electrodes, thereby recording the electrical activity during epileptic seizures. The EEG dataset encompasses both pre-ictal and ictal data, organized in a matrix format of channels over time, with a sampling rate of 256 points per second. Here, each electrode signal undergoes rescaling to an interval of $[-0.5,0.5]$ and is subsequently subsampled by averaging every sixteen points, resulting in the dataset with the time step $\Delta t=0.0625$. In Fig. 2, there is a sample image in which the doctor labels $Time=500$ as the separation time for the pre-ictal and ictal periods.
@@ 3.2 Dimension Reduction with Directed Anisotropic Diffusion Map
To construct the diffusion matrix, the data from all electrodes at each time point is viewed as a high-dimensional node in the diffusive graph. Notably, electrode signals display abnormal fluctuations preceding the onset of an epileptic seizure. Timely detection of abnormalities in electrical signals holds the potential to provide early warnings for epileptic seizures. We suggest identifying early warnings within a lower-dimensional space through the application of directed anisotropic diffusion maps.
@@
The diffusion matrix is constructed by the diffusion kernel and shows the transition among all high-dimensional nodes in the diffusive graph, which exhibits the transition probability from one point to another. To find the appropriate $\nabla_{x}f$ in a data driven way, we introduce the gradient flow in $\mathbf{R}^{n}$,
${\dot{x}=-\nabla_{x}f(x)}.$ (9)
Here, the gradient flow $\nabla_{x}f$ at $x_{i}=x_{t_{i}}$ is approximated by $\nabla_{x}f(x_{t_{i}})\approx-\frac{(x_{t_{i+1}}-x_{t_{i-1}})}{2\Delta t}$, which will be substituted into Eq.(1) for kernel calulation. Note the stochastic derivative is approximated similar to the finite difference formula, indicating the dynamic trend of stochastic process $x(t)$.
Figure 3: Comparison of diffusion map and directed anisotropic diffusion map. The first two columns (a) (b) represent the kernel matrices, and (c) (d) give the eigenvalues for diffusion map and directed anisotropic diffusion map respectively. For all the cases, we set $\epsilon=1$. Note that (a) (c) are constructed without the directed term, while (b) (d) are calculated with the data driven directed term learned by Eq.(9) .
@@
Then we can obtain the kernel matrices from the preprocessed dataset. As shown in Fig. 3, most points have a probability greater than $0.5$ with another point in $(a)$, indicating that the transition probability to other points is relatively equivalent. In $(b)$, the data points in the red frame have higher transition probabilities to each other, while the transition probabilities to the points outside gradually decrease. Through directed diffusion, the transition probability of sample points from inside the red box to the outside is reduced, and similar when points transit from outside to the inside. In this way, the diffusion intensity between points inside and outside the red box is forced to attenuate, so the two regions (’pre-ictal’ and ’ictal’ stages) could be better separated.
The diffusion matrix has a discrete sequence of eigenvalues $\{\lambda_{i}\}_{i\geq 1}$ and eigenvectors $\{\phi_{i}\}_{i\geq 1}$ such that $1\geq\lambda_{1}\geq\lambda_{2}\geq\cdots$. Note that for isotropic diffusion, the maximum eigenvalue is equal to 1, while for directed anisotropic diffusion, the largest eigenvalue is less than 1. Here, to illustrate the whole framework of our method in a straightforward way, we choose the spectral gap to help us find the effective dimension of latent space. In Fig. 3, (c) shows the first four eigenvalues of the diffusion map have similar gaps, while the first two eigenvalues of the directed anisotropic diffusion map in (d) have a bigger spectral gap than others. Thus we choose the latent space to be one-dimensional under directed anisotropic diffusion map. Considering some of the large eigenvalues are associated with the harmonics of earlier ones, there are more algorithms to find the effective embedding dimension [30], which can be further studied in the future.
@@
Figure 4: Comparison of low-dimensional embedding patterns with (a) diffusion map and (b) directed anisotropic diffusion map.
Now we compare the data projected onto the latent coordinate $\Phi_{1}(x)=\lambda_{1}\phi_{1}(x)$ with two different diffusion maps. In Fig. 4, both the projected data shows distribution shift at $Time\approx 600$. However, in $(b)$, the data after the ’shift’ has bigger variance and large deviated mean value compared with that before $Time\approx 600$. While in $(a)$, the data points before and after the shift have almost the same expectation, with bigger variance in the latter time period. We can see these two different patterns affect the dynamics of latent stochastic differential equations in the next section.
@@ 3.3 Latent Stochastic Dynamical Systems
In this section, we aim to elucidate the connection between the latent stochastic differential equation and the original high-dimensional time evolutionary data.
Following dimension reduction through diffusion map, we obtain the projected dataset onto the first diffusion coordinate $\Phi_{1}$. The projected data plotted over time displays the transition between two states: pre-ictal state and ictal state. Therefore, we assume that the stochastic differential equation in the latent space has two meta-stable states and rescale the data to an interval of $[-1,1]$ for convenience. Given data in Fig. 4 (b), we would like to obtain the latent stochastic differential equation in the form of Eq. (4). To get meaningful representations of drift and diffusion coefficients, we employ neural networks to approximate the latent equations. The inputs of networks are the polynomial basis functions of the coordinates and the time step is $0.0625$. The latent space contains $1500$ time points and $1499$ pairs of $\{(z(t_{i+1}),z(t_{i}))\}$ are available. We randomly select $1200$ pairs to train the networks and validate the identification accuracy on the rest $299$ pairs.
To find the explicit form of drift and diffusion coefficients in the unknown stochastic differential equations, all the unknown drift and diffusion terms are parameterized with neural networks. Then using the loglikelihood method in Sec. 2.2, the learnt stochastic differential equation under the coordinate $\Phi$ has the form
$dz(t)=\mu(z(t))dt+{\eta}(z(t))dB_{t},$ (10)
where for diffusion map, $\mu(z)=-0.0650+0.4767z+0.5048z^{2}-1.5105z^{3}$ and $\eta(z)=-0.4711$, and for directed diffusion map, $\mu(z)=0.1849+0.5058z-0.6381z^{2}-1.4880z^{3}$ and $\eta(z)=-0.4896$.
Moreover, we adopt density as the validation metric for the testing data (consisting of $299$ pairs). In Fig. 5, we compare the density of the estimated $z(t_{i+1})$ with the ground truth from corresponding projected data points in Fig.4. The alignment between our approximation and the true data density is evident. Additionally, the presence of two peaks in the density in Fig. 5 (b) signifies the existence of two meta-stable states in the projected coordinate by directed anisotropic diffusion map. However, both the densities in (a) only have the single peak.
Figure 5: The validation of the neural network approximation with regard to $\Phi_{1}$. (a) Diffusion map. (b) Directed anisotropic diffusion map.
@@ 3.4 Early Warning Indicators
For $\Phi_{1}$, we employ a time series comprising $1500$ data points to calculate the indicator over time and compare it with the standard deviation. The time window $l$ is set to be $300$, moving forward in each time unit $\Delta t=0.0625$. We propose the Onsager-Machlup indicator as the ratio of two Onsager-Machlup functionals (Eq. (6)) in each time window, equal to $\frac{OM(l,t)}{OM(0.9l,t-0.9l)}$. When applying the Eq. (7) to compute the sample entropy, the parameters are kept constant at $m=5$, $p=10$, $q=1$, and $r=2$. The values of the two indicators over time are red in Fig. 6 and the standard deviation is green.
First, considering the two patterns of projected data in Fig. 4, we can see that the standard deviations (green) in Fig. 6 with diffusion map (left column) and directed diffusion map (right column) have interpretable shapes. One interesting observation we would like to mention is that the directed diffusion map also helps the standard deviation(Std) indicator sharply change at $Time\approx 600$, which is quite different from the slowly increasing shape of Std under the diffusion map in the left column.
Next, in the top row of Fig. 6, both the Onsager-Machlup indicators with two kinds of diffusion maps show early warning signals before the standard deviation indicator booms up. The difference is that the directed diffusion map (b) makes the signal more narrow and peaky, while the diffusion map (a) makes the signal noisy and slowly changing at $Time\approx 600$. Furthermore, in the bottom of Fig. 6, both the sample entropy indicators with two types of diffusion maps present earlier warnings, compared with standard deviation. We can even find the benefit from directed diffusion map (d) that two more peaks of the entropy indicator ($Time\approx 900$ and $Time\approx 1200$) align well with the drop of the standard deviation. Consequently, the detection of change points can assist doctors in distinguishing pre-ictal and ictal periods, facilitating automated labeling processes.
Figure 6: The Onsager-Machlup indicator (OM) and the sample entropy (SE) of raw data projected onto coordinates $\Phi_{1}$ by diffusion map and the directed anisotropic diffusion map, compared with the standard deviation(Std)(green). (a) OM(Diffusion map). (b) OM(Directed diffusion map). (c) SE(Diffusion map). (d) SE(Directed diffusion map).
@@
Fig.7 shows the transition probability, assigning the width of each time window as $100$. The first $100$ points are selected as starting points, generating $100$ samples. For the directed anisotropic diffusion map, the split point is set as $-0.75$, delineating regions $A=(-\infty,-0.75]$ and $B=(-0.75,\infty)$. Similarly, the regions are $A=[-0.25,0.25]$ and $B=\mathbf{R}\backslash A$ for the diffusion map. In Fig. 7, we calculate the transition probability by Eq. (8) of state for each time step $t$ starting from the initial state. Assuming a transition occurs when the probability surpasses $0.5$, we designate the point with a probability equal to $0.5$ as the warning point. The warning points appear at $Time\approx 600$ for both two types of diffusion maps. However, the transition probability of the directed anisotropic diffusion map is much smoother, representing the two regions are better separated more robustly.
Figure 7: The transition probability of raw data projected onto coordinate $\Phi_{1}$. The red dot indicates the transition probability equal to $0.5$. (a) Diffusion map. (b) Directed anisotropic diffusion map.
3.5 Validation
We select the epileptic data of the same patient at a different visiting time to the hospital to serve as the validation set for our framework. To conduct out-of-sample verification for the indicators, we project new data points onto the embedded manifold by the geometric harmonics method [6, 31]. Specifically, the coordinates derived from the directed diffusion maps serve as a basis for the embedded manifold, and the formulation of the latent extension for the new point $x^{*}$ is carried out by $\Phi(x^{*})=[\Phi_{1}(x^{*})\ldots\Phi_{d}(x^{*})]$, where
${\Phi_{i}(x^{*})=\frac{1}{\lambda_{i}}\Sigma_{j}\frac{k_{\varepsilon}(x^{*},x_{j})}{\Sigma_{j}k_{\varepsilon}(x^{*},x_{j})}\Phi_{i}(x_{j}),\quad i=1,\ldots,d.}$
The validation set undergoes the same data preprocessing steps, and the new dataset (Fig. 8 (a)) comprises $1000$ data points, each with a time step of $0.0625$. The initial $500$ points represent pre-ictal data, followed by ictal data, labeled by medical professionals. The low-dimensional embedding of the new sample, projected onto the latent space, is shown in Fig. 8 (b). Moreover, three indicators are presented in (c), (d) and (e), with the same time window length $300$ and the latent stochastic differential equation learned in Section 3.3 under directed diffusion map. Notably, the upward trend of the Onsager-Machlup indicator outperforms that of the standard deviation in (c), and the sample entropy exhibits an abnormal spike before $Time=400$ in (d). Unfortunately, the transition probability indicator in (e) exhibits suboptimal performance. Despite the imperfect performance of one indicator, the new validation set underscores the feasibility of employing our method for the automated detection of abnormal signals in epileptic seizures, circumventing the need for manual diagnosis by medical professionals.
Figure 8: The out-of-sample points validate the effectiveness of our framework. (a) Validation dataset. (b) Embedding pattern. (c) OM indicator(red) compared with the Std(green). (d) SE indicator(red) compared with the Std(green). (e) The transition probability.
@@ 4 Conclusion
In this paper, we propose an efficient method for automatically identifying abnormalities in EEG signals. Our approach leverages directed anisotropic diffusion map for dimensionality reduction and establishes stochastic dynamical system in the low-dimensional latent space, which facilitates us to design early warning signals more efficiently. Besides, we also apply the framework to new dataset as validation.
@@
Early warning of epileptic seizures is of paramount importance for epileptic patients. The abrupt change is caught for early warning in the latent space, where normal state and ictal state can be viewed as two meta-stable states. The ability to identify transitions between meta-stable states plays a pivotal role in predicting and controlling brain behavior. We derive three effective warning signals—namely, the Onsager-Machlup indicator, the sample entropy indicator, and the transition probability indicator—utilizing information from the latent coordinates and the latent stochastic dynamical systems. These indicators enhance the robustness and accuracy of early warning systems for epileptic seizures. Furthermore, the computational cost of calculating these indicators from low dimensional data is much lower than that of the original high-dimensional data. This framework of learning latent stochastic systems and detecting abnormal dynamics has the potential to extend to general scenarios for other complex high dimensional time evolutionary data.